A study found that 28% of car owners in Fiji had their cars washed professionally rather than do it themselves. If 18 car owners are randomly selected, find the probability that at most two people have their cars washed professionally.


1
Expert's answer
2021-05-25T09:02:34-0400

Let "X=" the number of car owners who have their cars washed professionally: "X\sim Bin(n, p)."


"P(X=x)=\dbinom{n}{x}p^x(1-p)^{n-x}"

Given "p=0.28, n=18"


"P(X\leq2)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{18}{0}0.28^0(1-0.28)^{18-0}"

"+\dbinom{18}{1}0.28^1(1-0.28)^{18-1}"

"+\dbinom{18}{2}0.28^2(1-0.28)^{18-2}"

"=0.72^{18}+5.04\cdot0.72^{17}+11.9952\cdot0.72^{16}"

"\approx0.084195"